OpenStudy (12g):

A weather balloon holds 2,000 cubic meters of helium. The density of helium is 0.1765 kilograms per cubic meter. How many kilograms of helium does the balloon contain? Only enter a numerical value in the answer blank.

OpenStudy (12g):

Do you just do 2000 x 0.1765

OpenStudy (12g):

@Qwertty123

OpenStudy (12g):

@ganeshie8

OpenStudy (qwertty123):

Okay hold on!

OpenStudy (qwertty123):

Okay so d=m/v m=d*v v=d/m Where d denotes density, m denotes mass and v denotes volume by kc_kennylau

OpenStudy (anonymous):

(x/2000)=(.1765/1)

OpenStudy (12g):

so it is 353

OpenStudy (anonymous):

yes

OpenStudy (qwertty123):

YEAH! :D Great Job! :)

OpenStudy (12g):

that was right thanks so much

OpenStudy (qwertty123):

You Welcome! :D